OpenStudy (anonymous):

Prove sin(A±B)=sinAcosB ± cosAsinB using the following formulae: sinθ=cos(90º-θ) cos(A-B)=cosAcosB + sinAsinB cos(A+B)=cosAcosB - sinAsinB Hint: sin(A+B)=cos(90º-(A+B))

OpenStudy (shubhamsrg):

in cos(a-b) formulla,,substitute b as (90-b) ..[which means on both sides,ofcorse] also note that cos(-x)=cos x so you get the formulla for sin(a+b) now for sin(a-b),,make a similar subsn in cos(a+b) formulla.. are you getting it buddy ?

OpenStudy (anonymous):

Ummmm not quite sure. Would you be able to show me the first bit please?

OpenStudy (shubhamsrg):

hmm.. i'll show you 1st part..you show me 2nd after making subsn,,we have : cos(a-(90-b))=cosa cos(90-b) + sina sin(90-b) =>cos(a+b-90) = cos(90-(a+b))=cosa sinb + sina cosb = sin(a+b) hope that was fine.. so your turn now! ;)

OpenStudy (anonymous):

Thank you :)